Math 042: Exam II Date: April 02, 2010 Name: Section:

1. Which of the following is the graph of 3x + y = −6?

2. What is the slope of the line given by the equation x = −2?

(a) 1 (b) 0 (c) −2 (d) undefined

2  Exam II  April 02, 2010

3. Find the slope and the equation of the line that contains the point (3, −1) and is perpendicular to the line −2x − y = 4. Slope: . Equation: .

4. Find the equation of the line passing through (2, −5) and (−3, 3). Equation: .

5. Solve the following system of equations.

2x − y = 1
−3x + 2y = 5

Your answer: .

6. Graph 2x + 3y > 6.

7. Which of the following is an expansion of (4x − 3y)^{2}?

(a) 16x^{2} + 9y^{2} (b) 16x^{2} − 9y^{2} (c) 16x^{2} − 12xy + 9y^{2} (d) 16x^{2} − 24xy + 9y^{2}

8. Expand (x + 3)^{3}. Your answer: .

4  Exam II  April 02, 2010

9. y − 36y^{3} = .

10. Factor bx − by − 7x + 7y. Your answer: .

11. Simplify \frac{(2x^{−3}y)^{3}}{2x^{−2}y^{2}} . Your answer: .

12. Factor 2x^{2} − x − 10. Your answer: .

13. Solve x^{2} + 2x − 8 = 0. Your answer: .

14. Which one of the following is a factor of 8x^{2} − 40x?

(a) 8x^{2} (b) x^{2} − 5 (c) x − 5 (d) x^{2} − 5x

15. (\frac{3}{2})^{−3} = .

(a) −\frac{8}{27} (b) \frac{8}{27} (c) −\frac{27}{8} (d) \frac{27}{8}